OpenStudy (anonymous):

Solve for x. 4x(x-4)-5x(x-3)=0 x=

OpenStudy (anonymous):

x = -1, 0

OpenStudy (anonymous):

do you need more detail mizz?

myininaya (myininaya):

4x(x-4)-5x(x-3) = 4x^2-16x-5x^2+15x = -x^2-x = -x(x+1) = 0 =>-x=0 or x+1=0 =>x=0 or x=-1

OpenStudy (anonymous):

distribute the 4x and the -5x to get 4x^2 -16x -5x^2 +15x combine all like terms to get -x^2 - x factor out an x to get x(-x - 1) x = 0, and x = -1